OpenStudy (anonymous):

PLEASE HELP!!! For a simple random sample of forty 2005 Honda Accords, the mean gas mileage was 23 miles per gallon with a standard deviation of 1.5 miles per gallon. a) Construct a 95% confidence interval for the mean gas mileage for all 2005 Honda Accords. b) From the interval constructed in part (a), is there enough evidence to indicate that the typical 2005 Honda Accord has a gas mileage under 24 miles per gallon? Why or why not?

OpenStudy (amistre64):

\[\Large mean\pm z_{.025}\frac{sd}{\sqrt{n}}\]

OpenStudy (amistre64):

that rings a bell for some reason

OpenStudy (anonymous):

so would the setup be: \[23\pm1.96\frac{ 1.5 }{ \sqrt{40} }\]

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

for the second part, I ended up with \[22.53<\mu <23.47\]